# Birkhoff–Kellogg type results in product spaces and their application to differential systems

### v1

@@
Abstract. We provide a new version of the well-known Birkhoff-Kellogg invariant-direction Theorem in product spaces. Our results concern operator systems and give the existence of component-wise eigenvalues, instead of scalar eigenvalues as in the classical case, that have corresponding eigenvectors with all components nontrivial and localized by their norm. We also show that, when applied to nonlinear eigenvalue problems for differential equations, this localization property of the eigenvectors provides, in turn, qualitative properties of the solutions. This is illustrated in two context of systems of PDEs and ODEs. We illustrate the applicability of our theoretical results with two explicit examples.
2020 MSC: Primary 47H10, secondary 34B08, 35J57, 45G15, 47H11.
Key words and phrases. Birkhoff–Kellogg type result, nonlinear eigenvalue problem, operator system, nontrivial solution.
@@ 1. Introduction
A celebrated result in Nonlinear Analysis is the Birkhoff-Kellogg invariant-direction Theorem [5]. This theorem has been object of extensive research in the past and finds significant applications in the study of nonlinear eigenvalue problems in infinite-dimensional normed linear spaces, see for example the books [2, 12, 18], the recent papers [7, 15], and references therein. In the version by Krasnosel’skiĭ and Ladyženskiĭ [19], a similar result is set in cones of real Banach spaces, yielding the existence of a pair, constituted by a positive eigenvalue and an eigenvector, the latter localized inside a cone. We stress that a notable, common feature of the two above mentioned theorems is that they provide a localization of the eigenfunction; this localization, in turn, provides qualitative properties of the solution in the context of applications to nonlinear eigenvalue problems for differential equations. In the framework of systems the situation is somewhat more delicate. In fact, a direct application of one of the previous two results in product spaces would provide the existence of an eigenvalue with a corresponding (vectorial) eigenfunction that may have trivial components; this issue has been been highlighted for example in [15, Definition 1].
In this note we present a new version of the Birkhoff–Kellogg Theorem in product spaces, which, instead of a scalar eigenvalue, provides the existence of component-wise eigenvalues that have corresponding eigenvectors with all components nontrivial and localized by their norm. In the context of systems of integral equations and their applications, component-wise eigenvalues have been investigated in Chapter 3 of the book [1], where the authors sought constant-sign eigenvectors by means of topological tools such as the Schauder and the Krasnosel’skiĭ-Guo fixed point theorems, and in the book [14] where the authors sought, also via topological fixed point theory, the existence of positive eigenvectors.
Namely, we study systems of type
$\left\{\begin{array}[]{l}x=\lambda_{1}\,T_{1}(x,y),\\ y=\lambda_{2}\,T_{2}(x,y),\end{array}\right$. (1.1)
@@
In the context of fixed point theory, a component-wise approach has been utilized in the past, see for example [3, 4, 16, 17, 23, 24]; here we develop a somewhat analogue theory in the framework of nonlinear spectral theory. In particular, we give quite natural conditions yielding the existence of component-wise eigenvalues $\lambda_{1},\lambda_{2}>0$ and corresponding eigenfunctions $x_{0},y_{0}$, both of prescribed non-zero norm. We also provide further results in the settings (2) and (3), which yield the additional existence of negative eigenvalues with corresponding eigenfunctions.
Our existence results are motivated by the applications; in particular we show how our theory can be applied to differential systems. More in details, firstly we focus on the systems of PDEs
@@
where $\Omega\subset\mathbb{R}^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$ and $f,g$ are suitable continuous functions. Problems of this type are well-studied, both in the classical case of the spectral problem, that is with $\lambda=\lambda_{1}=\lambda_{2}$, see for example [6, 9, 16, 22], and in the case of component-wise eigenvalues with possibly different $\lambda_{1},\lambda_{2}$, see for instance [8, 13, 21] and references therein. In Theorem 3.1 below we provide sufficient conditions for (1.2) that ensure the existence of a pair of positive component-wise eigenvalues with eigenfunctions possessing nontrivial components with localized norms. The result is then illustrated in a specific example.
Our second setting of application is a BVP for the system of ODEs,
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}f(t,u(t),v(t))=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}g(t,u(t),v(t))=0,\ t\in[0,1],\\ u(0)=u(1)=0=v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1),\end{array}\right$. (1.3)
where $f,g$ are suitable continuous functions. The BCs that occur in (1.3) have been investigated for a different set of parameters by Lan[20]. We rewrite the system (1.3) in terms of a system of Hammerstein integral equations, and we work in the Cartesian product of a conical shell times a ball in the space of continuous functions. Also in this case, our approach yields the existence of two distinct pairs of component-wise eigenvalues with nontrivial eigenfunctions; this is illustrated in a toy model as well.
Overall results are new from both the theoretical and the applied point of view and complement the ones in [1, 2, 6, 8, 9, 13, 14, 15, 16, 21, 22, 24].
@@ 2. Birkhoff-Kellogg type results
We begin this Section by recalling the classical Birkhoff-Kellogg invariant-direction Theorem [5], cf. [11, Theorem 6.1].
@@
In the following version by Krasnosel’skiĭ and Ladyženskiĭ [19], cf. [18, Theorem 5.5], a similar result is set in cones of real Banach spaces; we recall that a cone $K$ of a normed linear space $(X,\|\,\|)$ is a closed set with $K+K\subset K$, $\mu K\subset K$ for all $\mu\geq 0$ and $K\cap(-K)=\{0\}$.
@@
With abuse of notation (the whole space $X$ is not a cone) we will still denote $X_{r}=B_{r}$, $\overline{X}_{r}=\overline{B}_{r}$, $\partial\overline{X}_{r}=\partial\overline{B}_{r}$, so that, if $X$ is infinite dimensional, again $\partial\overline{X}_{r}$ is a retract of $\overline{X}_{r}$.
@@
The following result is a version of the Birkhoff–Kellogg Theorem in product spaces.
@@
$N(x,y)=\left(r_{1}\dfrac{T_{1}\left(\rho_{1}(x),y\right)}{\left\|T_{1}\left(\rho_{1}(x),y\right)\right\|},r_{2}\dfrac{T_{2}\left(x,\rho_{2}(y)\right)}{\left\|T_{2}\left(x,\rho_{2}(y)\right)\right\|}\right),$
and observe that, by (2.1), $N$ is well-defined. Since $N$ is a compact map, Schauder fixed point theorem (see e.g. [11, §6, Theorem 3.2]) ensures that $N$ has at least one fixed point $(x_{0},y_{0})\in\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}$. Observe that $N\left(\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}\right)\subset\partial\overline{C}_{1,r_{1}}\times\partial\overline{C}_{2,r_{2}}$ and so it follows that $(x_{0},y_{0})\in\partial\overline{C}_{1,r_{1}}\times\partial\overline{C}_{2,r_{2}}$, that is,
@@
$\left\{\begin{array}[]{l}x_{0}=\lambda_{0}\,T_{1}(x_{0},y_{0}),\\ y_{0}=\lambda_{0}\,T_{2}(x_{0},y_{0})\end{array}\right$. (2.3)
cannot be guaranteed. Indeed, consider as Banach spaces $X=Y=\mathbb{R}$, the cones $K_{1}=K_{2}=[0,+\infty)$, $r_{1}=r_{2}=1$ and the continuous function $T:[0,1]\times[0,1]\rightarrow[0,+\infty)\times[0,+\infty)$ given by
@@
Note that $\inf_{x=1}\left|T_{1}(x,y)\right|=2>0$ and $\inf_{y=1}\left|T_{2}(x,y)\right|=3>0$, but there is no $\lambda\in(0,+\infty)$ such that $(1,1)=\lambda\,T(1,1)=\lambda\,(3,4)$.
@@
Let $K_{1}$ be a cone in the normed linear space $X_{1}$ and $X_{2}$ be an infinite dimensional normed space. Let $r_{1},r_{2}$ be positive constants and suppose that $T=(T_{1},T_{2}):\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}\longrightarrow K_{1}\times X_{2}$ is a compact map satisfying that
$\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|\leq r_{2}}\left\|T_{1}(x,y)\right\|>0\quad\text{and}\quad\inf_{\left\|x\right\|\leq r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{2}(x,y)\right\|>0.$
Then there exist $\lambda_{1,1},\lambda_{2,1},\lambda_{1,2}>0$, $\lambda_{2,2}<0$ and $(x_{0,j},y_{0,j})\in\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}$, $j=1,2$, with $\left\|x_{0,j}\right\|=r_{1}$ and $\left\|y_{0,j}\right\|=r_{2}$ such that
@@
The first solution is ensured by Theorem 2.3. In order to obtain the second one, just consider the map $\tilde{N}:\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}\rightarrow\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}$ given by
$\tilde{N}(x,y)=\left(r_{1}\dfrac{T_{1}\left(\rho_{1}(x),y\right)}{\left\|T_{1}\left(\rho_{1}(x),y\right)\right\|},-r_{2}\dfrac{T_{2}\left(x,\rho_{2}(y)\right)}{\left\|T_{2}\left(x,\rho_{2}(y)\right)\right\|}\right).$
As a consequence of Schauder fixed point theorem, $\tilde{N}$ has a fixed point $(x_{0,2},y_{0,2})$ located in $\partial\overline{K}_{1,r_{1}}\times\partial\overline{B}_{r_{2}}$, that is,
@@
Under the assumptions of Theorem 2.3, if both $C_{1}$ and $C_{2}$ are infinite dimensional normed spaces, then there exist four couples of numbers $\lambda_{1},\lambda_{2}$ and associated points $(x_{0},y_{0})\in\overline{B}_{r_{1}}\times\overline{B}_{r_{2}}$ with $\left\|x_{0}\right\|=r_{1}$ and $\left\|y_{0}\right\|=r_{2}$ such that (2.2) holds. Indeed, it suffices to apply the Schauder theorem to each auxiliary map
$N_{j,k}(x,y)=\left((-1)^{j}r_{1}\dfrac{T_{1}\left(\rho_{1}(x),y\right)}{\left\|T_{1}\left(\rho_{1}(x),y\right)\right\|},(-1)^{k}r_{2}\dfrac{T_{2}\left(x,\rho_{2}(y)\right)}{\left\|T_{2}\left(x,\rho_{2}(y)\right)\right\|}\right),\quad j,k\in\{1,2\}.$

### v2

@@
Abstract.
We provide a new version of the well-known Birkhoff-Kellogg invariant-direction Theorem in product spaces. Our results concern operator systems and give the existence of component-wise eigenvalues, instead of scalar eigenvalues as in the classical case, that have corresponding eigenvectors with all components nontrivial and localized by their norm. We also show that, when applied to nonlinear eigenvalue problems for differential equations, this localization property of the eigenvectors provides, in turn, qualitative properties of the solutions. This is illustrated in two contexts of systems of PDEs and ODEs. We show the applicability of our theoretical results with two explicit examples.
Key words and phrases:
Birkhoff–Kellogg type result, nonlinear eigenvalue problem, operator system, nontrivial solution
2020 Mathematics Subject Classification:
Primary 47H10, secondary 34B08, 35J57, 45G15, 47H11
@@ 1. Introduction
A celebrated result in Nonlinear Analysis is the Birkhoff-Kellogg invariant-direction Theorem [5]. This theorem has been object of extensive research in the past and finds significant applications in the study of nonlinear eigenvalue problems in infinite-dimensional normed linear spaces, see for example the books [2, 13, 19], the recent papers [7, 16], and references therein. In the version by Krasnosel’skiĭ and Ladyženskiĭ [20], a similar result is set in cones of real Banach spaces, yielding the existence of a pair, constituted by a positive eigenvalue and an eigenvector, the latter localized inside a cone. We stress that a notable, common feature of the two above mentioned theorems is that they provide a localization of the eigenfunction; this localization, in turn, provides qualitative properties of the solution in the context of applications to nonlinear eigenvalue problems for differential equations. In the framework of systems the situation is somewhat more delicate. In fact, a direct application of one of the previous two results in product spaces would provide the existence of an eigenvalue with a corresponding (vectorial) eigenfunction that may have trivial components; this issue has been been highlighted for example in [16, Definition 1].
In this paper we present a new version of the Birkhoff–Kellogg Theorem in product spaces, which, instead of a scalar eigenvalue, provides the existence of component-wise eigenvalues that have corresponding eigenvectors with all components nontrivial and localized by their norm. In the context of systems of integral equations and their applications, component-wise eigenvalues have been investigated in Chapter 3 of the book [1], where the authors sought constant-sign eigenvectors by means of topological tools such as the Schauder and the Krasnosel’skiĭ-Guo fixed point theorems, and in the book [15] where the authors sought, also via topological fixed point theory, the existence of positive eigenvectors. More in details, we study systems of type
$\left\{\begin{array}[]{l}x=\lambda_{1}\,T_{1}(x,y),\\ y=\lambda_{2}\,T_{2}(x,y),\end{array}\right$.
@@
In the context of fixed point theory, a component-wise approach has been utilized in the past, see for example [3, 4, 17, 18, 27, 28]; here we develop a somewhat analogue theory in the framework of nonlinear spectral theory. In particular, we give quite natural conditions yielding the existence of component-wise eigenvalues $\lambda_{1},\lambda_{2}>0$ and corresponding eigenfunctions $x_{0},y_{0}$, both of prescribed non-zero norm. We also provide further results in the settings (2) and (3), which yield the additional existence of negative eigenvalues with corresponding eigenfunctions. Furthermore we present a direct application of the classical result by Krasnosel’skiĭ and Ladyženskiĭ to the context of systems and present a comparison between the two approaches, see Remark 2.8.
Section 3 of the paper is devoted to applications of the theoretical results to differential systems. We stress that these applications are given as a motivation for the theoretical results and are provided for illustrative purposes. In order to keep the paper as readable as possible, we study these problems in the case of continuous nonlinearities.
We firstly focus on the systems of PDEs
@@
where $\Omega\subset\mathbb{R}^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$ and $f,g$ are suitable non-negative continuous functions. Problems of this type are well-studied, both in the classical case of the spectral problem, that is with $\lambda=\lambda_{1}=\lambda_{2}$, see for example [6, 9, 17, 26], and in the case of component-wise eigenvalues with possibly different $\lambda_{1},\lambda_{2}$, see for instance [8, 14, 24, 25] and references therein. Within this framework, we utilize the cone product setting (1); in Theorem 3.1 we provide sufficient conditions for (1.1) that ensure the existence of a pair of positive component-wise eigenvalues with eigenfunctions possessing nontrivial components with localized norms. The result is then illustrated in a specific example.
Our second setting of application is the following BVP for a system of ODEs
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}f(t,u(t),v(t))=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}g(t,u(t),v(t))=0,\ t\in[0,1],\\ u^{\prime}(0)=u(1)+u^{\prime}(1)=0,\\ v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1)=0.\end{array}\right$. (1.2)
where $f,g$ are suitable non-negative continuous functions. The BCs that occur in (1.2) have been investigated for a different set of parameters by Lan [21, 25]. We rewrite the system (1.2) in terms of a system of Hammerstein integral equations. Since one of the two corresponding kernels changes sign, we use the setting (2), namely the Cartesian product of a conical shell times a ball in the space of continuous functions. In this case our approach yields the existence of two distinct pairs of component-wise eigenvalues with nontrivial eigenfunctions possessing nontrivial components; this is illustrated in a toy model as well.
Overall results are new, from both the theoretical and the applied point of view, and complement the ones in [1, 2, 6, 8, 9, 14, 15, 16, 17, 24, 25, 26, 28].
@@ 2. Birkhoff-Kellogg type results
We begin this Section by recalling the classical Birkhoff-Kellogg invariant-direction Theorem [5], cf. [12, Theorem 6.1].
@@
In the following version by Krasnosel’skiĭ and Ladyženskiĭ [20], cf. [19, Theorem 5.5], a similar result is set in cones of real Banach spaces; we recall that a cone $K$ of a normed linear space $(X,\|\,\|)$ is a closed set with $K+K\subset K$, $\mu K\subset K$ for all $\mu\geq 0$ and $K\cap(-K)=\{0\}$.
@@
With abuse of notation (the whole space $X$, when nontrivial, is not a cone) we will still denote $X_{r}=B_{r}$, $\overline{X}_{r}=\overline{B}_{r}$, $\partial\overline{X}_{r}=\partial\overline{B}_{r}$, so that, if $X$ is infinite dimensional, again $\partial\overline{X}_{r}$ is a retract of $\overline{X}_{r}$.
@@
In the following we will consider the product space $X_{1}\times X_{2}$ endowed with the maximum norm. With abuse of notation, the norm in any space will be still denoted by $\|\ \|$. The next result is a version of the Birkhoff–Kellogg Theorem in product spaces.
@@
$\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{1}(x,y)\right\|>0\quad\text{and}\quad\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{2}(x,y)\right\|>0.$ (2.1)
@@
$N(x,y)=\left(r_{1}\dfrac{T_{1}\left(\rho_{1}(x),\rho_{2}(y)\right)}{\left\|T_{1}\left(\rho_{1}(x),\rho_{2}(y)\right)\right\|},r_{2}\dfrac{T_{2}\left(\rho_{1}(x),\rho_{2}(y)\right)}{\left\|T_{2}\left(\rho_{1}(x),\rho_{2}(y)\right)\right\|}\right),$
and observe that, by (2.1), $N$ is well-defined. Since $N$ is a compact map, Schauder fixed point theorem (see e.g. [12, §6, Theorem 3.2]) ensures that $N$ has at least one fixed point $(x_{0},y_{0})\in\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}$. Observe that $N\left(\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}\right)\subset\partial\overline{C}_{1,r_{1}}\times\partial\overline{C}_{2,r_{2}}$ and so it follows that $(x_{0},y_{0})\in\partial\overline{C}_{1,r_{1}}\times\partial\overline{C}_{2,r_{2}}$, that is,
@@
$\left\{\begin{array}[]{l}x_{0}=\lambda_{0}\,T_{1}(x_{0},y_{0}),\\ y_{0}=\lambda_{0}\,T_{2}(x_{0},y_{0})\end{array}\right$.
cannot be guaranteed. Indeed, consider as Banach spaces $X_{1}=X_{2}=\mathbb{R}$, the cones $K_{1}=K_{2}=[0,+\infty)$, $r_{1}=r_{2}=1$ and the continuous function $T:[0,1]\times[0,1]\rightarrow[0,+\infty)\times[0,+\infty)$ given by
@@
Note that $\inf_{{\left\|x\right\|=r_{1},\ \left\|y\right\|=r_{2}}}\left|T_{1}(x,y)\right|=3>0$ and $\inf_{{\left\|x\right\|=r_{1},\ \left\|y\right\|=r_{2}}}\left|T_{2}(x,y)\right|=4>0$, but there is no $\lambda\in(0,+\infty)$ such that $(1,1)=\lambda\,T(1,1)=\lambda\,(3,4)$.
@@
Let $K_{1}$ be a cone in the normed linear space $X_{1}$, and $X_{2}$ be an infinite dimensional normed space. Let $r_{1},r_{2}$ be positive constants and suppose that $T=(T_{1},T_{2}):\overline{K}_{1,r_{1}}\times\overline{B}_{2,r_{2}}\longrightarrow K_{1}\times X_{2}$ is a compact map satisfying that
$\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{1}(x,y)\right\|>0\quad\text{and}\quad\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{2}(x,y)\right\|>0.$
Then there exist $\lambda_{1,1},\lambda_{2,1},\lambda_{1,2}>0$, $\lambda_{2,2}<0$ and $(x_{0,j},y_{0,j})\in\overline{K}_{1,r_{1}}\times\overline{B}_{2,r_{2}}$, $j=1,2$, with $\left\|x_{0,j}\right\|=r_{1}$ and $\left\|y_{0,j}\right\|=r_{2}$ such that
@@
The first solution is ensured by Theorem 2.3. In order to obtain the second one, just consider the map $\tilde{N}:\overline{K}_{1,r_{1}}\times\overline{B}_{2,r_{2}}\rightarrow\overline{K}_{1,r_{1}}\times\overline{B}_{2,r_{2}}$ given by
$\tilde{N}(x,y)=\left(r_{1}\dfrac{T_{1}\left(\rho_{1}(x),\rho_{2}(y)\right)}{\left\|T_{1}\left(\rho_{1}(x),\rho_{2}(y)\right)\right\|},-r_{2}\dfrac{T_{2}\left(\rho_{1}(x),\rho_{2}(y)\right)}{\left\|T_{2}\left(\rho_{1}(x),\rho_{2}(y)\right)\right\|}\right).$
As a consequence of Schauder fixed point theorem, $\tilde{N}$ has a fixed point $(x_{0,2},y_{0,2})$ located in $\partial\overline{K}_{1,r_{1}}\times\partial\overline{B}_{2,r_{2}}$, that is,
@@
Under the assumptions of Theorem 2.3, if both $C_{1}$ and $C_{2}$ are infinite dimensional normed spaces, then there exist four couples of numbers $\lambda_{1},\lambda_{2}$ and associated points $(x_{0},y_{0})\in\overline{B}_{1,r_{1}}\times\overline{B}_{2,r_{2}}$ with $\left\|x_{0}\right\|=r_{1}$ and $\left\|y_{0}\right\|=r_{2}$ such that (2.2) holds. Indeed, it suffices to apply the Schauder theorem to each auxiliary map
$N_{j,k}(x,y)=\left((-1)^{j}r_{1}\dfrac{T_{1}\left(\rho_{1}(x),\rho_{2}(y)\right)}{\left\|T_{1}\left(\rho_{1}(x),\rho_{2}(y)\right)\right\|},(-1)^{k}r_{2}\dfrac{T_{2}\left(\rho_{1}(x),\rho_{2}(y)\right)}{\left\|T_{2}\left(\rho_{1}(x),\rho_{2}(y)\right)\right\|}\right),\quad j,k\in\{1,2\}.$
Let us now stress, once again, that a key feature of our results is the localization of the solutions. In order to compare our results with the classical ones we provide now a direct application of Theorem 2.2 to the situation in which, for each $i\in\{1,2\}$, $C_{i}=K_{i}$ is a cone. An analogous statement holds when $C_{i}$, $i\in\{1,2\}$, is either a cone or an infinite dimensional normed space.
Theorem 2.7.
Let $X_{1}$ and $X_{2}$ be real Banach Spaces and let $K_{1}\subset X_{1}$, $K_{2}\subset X_{2}$ cones. Let $r_{1},r_{2}$ be positive constants and suppose that
$T=(T_{1},T_{2}):\overline{K}_{1,r_{1}}\times\overline{K}_{2,r_{2}}\longrightarrow K_{1}\times K_{2},$
is a compact map satisfying that
Then there exists $\lambda_{0}>0$ and $(x_{0},y_{0})\in\overline{K}_{1,r_{1}}\times\overline{K}_{2,r_{2}}$ with $\left\|x_{0}\right\|=r_{1}$ or $\left\|y_{0}\right\|=r_{2}$ such that
$\left\{\begin{array}[]{l}x_{0}=\lambda_{0}\,T_{1}(x_{0},y_{0}),\\ y_{0}=\lambda_{0}\,T_{2}(x_{0},y_{0}).\end{array}\right$. (2.4)
Proof.
First of all note that $K=K_{1}\times K_{2}$ is a cone in the space $X=X_{1}\times X_{2}$. Let $U={B}_{1,r_{1}}\times{B}_{2,r_{2}}$ and take $(x,y)\in K\cap\partial U$. Now either $(x,y)\in\partial B_{1,r_{1}}\times\overline{B}_{2,r_{2}}$ or $(x,y)\in\overline{B}_{1,r_{1}}\times\partial B_{2,r_{2}}$. In the first case we have
$\|T(x,y)\|\geq\|T_{1}(x,y)\|$
while in the second case we obtain
$\|T(x,y)\|\geq\|T_{2}(x,y)\|.$
In both cases, by condition (2.3), we have
$\|T(x,y)\|\geq\min\{\|T_{1}(x,y)\|,\|T_{2}(x,y)\|\}\\ \geq\min\Bigl\{\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|\leq r_{2}}\left\|T_{1}(x,y)\right\|,\inf_{\left\|x\right\|\leq r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{2}(x,y)\right\|\Bigr\}>0.$
This implies the inequality
$\inf_{(x,y)\in K\cap\partial U}\|T(x,y)\|>0.$
[…]
Remark 2.8.
Note that condition (2.3) is stronger than (2.1); this situation is particularly meaningful in the applications, see the hypothesis $b)$ in Theorem 3.3. Furthermore, note that Theorem 2.3 provides a better localization of the eigenfunctions (both of them are nontrivial), since Theorem 2.7 provides couples $(x_{0},y_{0})$ in which one of the components can be trivial.